# Doppler Correspondence: Non-Iterative Scan Matching With Doppler Velocity-Based Correspondence

### v2

@@ I Introduction
State estimation is a fundamental component in autonomous systems, enabling robots to achieve tasks accurately and safely [3, 7, 20]. Accurate odometry plays a critical role in reliable state estimation by tracking the robot’s movement and position. Therefore, LiDAR odometry has been extensively researched to achieve high accuracy and real-time performance, leveraging its superior precision and detailed environmental sensing capabilities. However, scan matching, which is essential for LiDAR odometry, is degraded in adverse weather conditions such as rain, smoke, and fog [15]. Additionally, most scan matching methods, which depend on the geometry of the environment, often fail to accurately estimate sensor motion in repetitive geometric structures such as tunnels and highways [7].
Recently, the frequency-modulated continuous wave (FMCW) 4D LiDAR and 4D radar have shown great potential in alleviating these limitations. Unlike conventional LiDAR, FMCW-based 4D ranging sensors are more resilient to weather conditions [10, 15]. Moreover, they provide 4D point cloud information (range, azimuth, elevation, and Doppler velocity), which has potential [7, 15] for robust scan matching in scenarios with repetitive geometric structures.
Existing approaches related to correspondence estimation for 4D point cloud scan matching either rely on heuristic utilization of radar cross section (RCS) [8] or leverage the uncertainty inherent in point clouds derived from 3D information [20]. Although FMCW-based 4D ranging sensors provide Doppler velocity but still the correspondence estimation methods typically rely solely on 3D information [6, 7, 11, 13, 14, 23].
@@
Figure 2: The result of the proposed non-iterative scan matching method using the nyl trajectory in the NTU4DRadLM dataset [22].
@@ II Related Work
4D radar odometry and simultaneous localization and mapping (SLAM) methods have gained significant attention due to their robustness in challenging environments, such as visually degraded or bad weather conditions. Early work by Doer and Trommer [3, 4] demonstrated the feasibility of fusing 4D radar data with inertial measurements to achieve 4D radar-inertial odometry (RIO). Using extended kalman filter (EKF)-based approaches, they highlighted improvements in pose estimation through techniques such as RANSAC-based ego-velocity estimation, barometric height fusion, and online 4D radar extrinsic calibration, which removed the need for tedious pre-calibration processes. Building on this foundation, Michalczyk et al. [13, 14] extended RIO methods, incorporating persistent landmark detection and multi-state estimation frameworks to improve performance.
More recently, as 4D radar point clouds become more dense, researchers have focused on optimization-based approaches. Unlike earlier Kalman filter-based methods, which rely on sequential state estimation, optimization-based methods aim to solve the sub-trajectory by considering the global consistency of sensor data. Li et al. [11] developed a pose graph optimization-based 4D radar SLAM framework that utilizes ego-velocity pre-integration factors to improve robustness in noisy 4D radar data. Huang et al. [8] introduced an enhanced RIO that integrates Doppler velocity and RCS information to filter noisy point cloud, adhering to the less is more principle. Additionally, Nissov et al. [15] tackled the challenges of LiDAR degeneracy by proposing a LiDAR-4D radar-inertial fusion, leveraging 4D radar’s robustness in the fog-filled hallway, while preserving LiDAR accuracy in well-conditioned environments. Herraez et al. [6] presented methods for pose estimation and map creation, highlighting the benefits of velocity-aided odometry and map filtering to enhance accuracy. Zhang et al. [23] and Xu et al. [20] further advanced 4D radar odometry by modeling point uncertainty in polar coordinates, integrating into data association and motion estimation for improved performance in adverse conditions.
@@
4D LiDAR is an emerging technology that enables Doppler velocity measurements, offering advantages in environments with geometric degeneracy or dynamic objects. Several recent studies have explored the potential of this novel sensing modality for odometry and SLAM tasks. Wu et al. [19] introduced the first continuous-time 4D LiDAR-only odometry method leveraging Doppler velocity measurements. Their method employs Gaussian process regression to estimate vehicle trajectories and correct motion distortion caused by scanning, significantly outperforming existing methods. Similarly, Hexsel et al. [7] proposed DICP, which integrates Doppler velocity measurements into the ICP framework for robust point cloud registration. By jointly optimizing Doppler and geometric objective functions, their method demonstrated improved registration accuracy and convergence speed, particularly in featureless environments such as tunnels and hallways. Yoon et al. [21] and Lisus et al. [12] developed correspondence-free methods that estimate 6-DOF velocity directly from Doppler information without requiring explicit scan matching and incorporate multiple sensors or complementary modules such as IMUs or gyroscopes for full motion estimation.
@@ II-C Summary and Limitations of 4D Radar / LiDAR Odometry
The 4D information of FMCW 4D LiDAR and 4D radar has leveraged its advantages in three main directions. First, Doppler velocity has been effectively incorporated into optimization frameworks to improve accuracy and robustness in challenging environments, such as repetitive geometric settings, foggy conditions, or a significant number of dynamic objects [3, 4, 6, 7, 11, 13, 14, 15, 19, 23]. Second, scan matching methods have been developed by refining point correspondences, using RCS information [8] or the uncertainty of point clouds [20]. Third, direct motion estimation methods [21, 12] attempt to recover full 6-DOF motion from Doppler information without relying on point correspondences.
@@ III-A Problem Definition
Mostly, the Correspondence Estimation step solely relies on the spatial information of each point cloud. Relying only on the spatial information in Correspondence Estimation makes registration prone to degradation in environments with repetitive geometric structures [7, 15] or when consecutive scans fail to capture the same spatial points due to its sparse and noisy characteristics [8]. To address these limitations, this study proposes Doppler Correspondence for the Correspondence Estimation, which is explained from the following in detail.
@@ III-B Geometric Derivation of Doppler Correspondence
Note that the above relationship, termed as Doppler Correspondence, is formulated such that the left-hand side, $f(p_{i},v_{p,i})$, depends only on the information from point $p_{i}$, while the right-hand side, $g(q_{j},v_{q,j})$, depends only on the information from point $q_{j}$. Furthermore, since there are no terms involving the $\Delta t$, this Doppler Correspondence is inherently invariant to translation of the sensor. Moreover, unlike traditional correspondence that relies only on spatial $(x,y,z)$ information, this method additionally utilizes Doppler information. This enables a more robust and reliable matching process in repetitive geometric environments, as it relies more on the flow of the point cloud [2] rather than the spatial information from the environments.
@@ III-C Kinematic Derivation of Doppler Correspondence
Alternative to the previous geometric derivation of Doppler Correspondence, this section utilizes manipulation of simple matrices and vectors. To establish Doppler Correspondence, we start with the kinematic relationship from  [2, 25]:
@@ III-D Outlier Rejection
This occurs because $g(q_{j},v_{q,j})$ is identical to $g(q^{\prime}_{j},v_{q^{\prime},j})$, as the range and radial velocity of each point is the same ($r_{q_{j}}=r_{q^{\prime}_{j}}\text{ and }v_{q_{j}}=v_{q^{\prime}_{j}}$). Such ambiguity can result in incorrect correspondence estimation. To address this issue, it is assumed that correctly matched points $p_{i}$ and $q_{j}$ lie within similar spatial regions between scans. This assumption enables the use of a distance-based rejection mechanism to filter out mismatched candidates, as shown in Fig. 4(b). Similar to the maximum correspondence distance used in [17], a threshold is applied to eliminate unlikely matches. Specifically, both the spatial distance $d_{\text{spatial}}=\|p_{i}-q_{j}\|_{2}$ and the Doppler distance $d_{\text{doppler}}=|f(p_{i},v_{p,i})-g(q_{j},v_{q,j})|$ should satisfy:
@@ III-E Non-iterative Scan Matching Algorithm
The process begins by receiving two consecutive scans, $\mathcal{P}$ and $\mathcal{Q}$, and calculating $f(p_{i},v_{p,i})\text{ and }g(q_{j},v_{q,j})$. FindClosest is then performed to estimate correspondences. Unlike the conventional ICP explained in Algorithm 2, which relies on 3D-based correspondence, our method operates on 1D-based correspondence. This could reduce computational complexity, even though $f(p_{i},v_{p,i})$ and $g(q_{j},v_{q,j})$ need to be calculated. While this approach may be slightly less accurate than 3D-based methods due to less information, it offers faster computations. Furthermore, unlike the closest point correspondence methods [1, 7, 16, 17], which iteratively refine correspondences through multiple iterations, Algorithm 1 performs a direct correspondence estimation in a single step. This eliminates the iterative search process, significantly reducing the computational load, as illustrated in Fig. 1. Next section, the feasibility of Doppler Correspondence is validated by comparing the proposed method against ICP-based approaches.
@@ IV-A1 Sparse 4D Radar
The Campus dataset [11] is utilized to evaluate sparse 4D radar point clouds. The data were collected using a ZF FRGen21 4D radar sensor mounted on the front bumper of a vehicle. Each 4D radar scan generates a sparse and noisy point cloud, containing approximately 400 to 1,400 points every 60 ms. The Campus dataset encompasses outdoor environments featuring trees, vehicles, and pedestrians. These diverse settings challenge methods to operate effectively under low-density radar point cloud conditions.
@@ IV-A2 Semi-Dense 4D Radar
The Loop 1 dataset from the NTU4DRadLM dataset [22] is used to evaluate semi-dense 4D radar. The data were collected using a car platform with an Oculii Eagle 4D radar, traversing the NTU campus main roads, resulting in a 6.95 km trajectory. The car moved at an average speed of 25–30 km/h, and the 4D radar scans exhibit semi-dense point clouds with approximately 1,000 to 2,000 points per 80ms. The environment features a mix of structured and semi-structured regions.
@@ IV-A3 Dense 4D LiDAR
The Straight Wall, Curved Wall, and Baker-Barry Tunnel datasets from [7] are utilized to evaluate dense 4D LiDAR. The Straight Wall and Curved Wall datasets, generated in the CARLA simulator, simulate featureless environments by placing large parallel walls along the trajectory. These datasets provide an effective platform for evaluating odometry performance, particularly in environments with limited geometric features and repetitive structures. The Baker-Barry Tunnel dataset was collected using an Aeva Aeries I FMCW 4D LiDAR sensor. The point clouds from 4D LiDAR are significantly denser than 4D radar data, containing approximately 40,000 to 120,000 points per 100ms, depending on the sequence.
@@ IV-B Experiment Setup
The proposed method is evaluated by comparing its performance with established methods: point-to-point ICP and DICP [7]. DICP is a novel algorithm that integrates Doppler velocity into the Optimization in the ICP framework. It enhances scan matching by jointly optimizing both the closest point term and additional constraints derived from the Doppler velocity of each point. These methods are tested in two conditions: No Seed and With Seed. In the No Seed category, no initial pose estimate is provided, whereas the With Seed category uses the estimated pose from the previous scan pair under a constant-velocity motion model assumption [18].
@@
[…]
@@
[…]
@@
[…]
@@
[…]
@@ IV-C Evaluation Metrics
The performance is evaluated using the relative error (RE) from [24]. This metric standardizes trajectory estimation accuracy, allowing fair comparisons across different methods. The evaluation measures odometry accuracy through segment-wise error analysis. This approach provides detailed insights into the local and global performance of the trajectory estimation. By using shorter sub-trajectories, RE captures local consistency and short-term accuracy, while longer sub-trajectories emphasize global consistency and long-term accuracy. This dual capability makes RE a versatile and informative metric for trajectory analysis.
@@ IV-D Experiment Results
The performance evaluation on sparse 4D radar data compares the proposed method against the standard point-to-point ICP approach, while the open-source implementation of DICP is excluded due to its failure to produce reliable results on sparse 4D radar data. Under these conditions, closest point correspondence, which relies on the spatial overlap of consecutive scans, becomes unreliable [8]. By contrast, the proposed correspondence leverages the flow of points [2] rather than explicit spatial positions, offering robustness against sparse and noisy data. As shown in Table I, Figs. 5(a) and (b), and Fig. 6(a), the proposed method attains substantially higher accuracy than that of point-to-point ICP. Moreover, unlike point-to-point ICP, which requires repeated correspondence estimation in 3D space, our approach simplifies the process to a single-step and 1D correspondence estimation.

### v3

@@ I Introduction
State estimation is a fundamental component in autonomous systems, enabling robots to achieve tasks accurately and safely [doer2020ekf, hexsel2022dicp, xu2024modeling]. Accurate odometry plays a critical role in reliable state estimation by tracking the robot’s movement and position. Therefore, LiDAR odometry has been extensively researched to achieve high accuracy and real-time performance, leveraging its superior precision and detailed environmental sensing capabilities. However, scan matching, which is essential for LiDAR odometry, is degraded in adverse weather conditions such as rain, smoke, and fog [nissov2024degradation]. Additionally, most scan matching methods, which depend on the geometry of the environment, often fail to accurately estimate sensor motion in repetitive geometric structures such as tunnels and highways [hexsel2022dicp].
Recently, the frequency-modulated continuous wave (FMCW) 4D LiDAR and 4D radar have shown great potential in alleviating these limitations. Unlike conventional LiDAR, FMCW-based 4D ranging sensors are more resilient to weather conditions [lee2023frequency, nissov2024degradation]. Moreover, they provide 4D point cloud information (range, azimuth, elevation, and Doppler velocity), which has potential [hexsel2022dicp, nissov2024degradation] for robust scan matching in scenarios with repetitive geometric structures.
Existing approaches related to correspondence estimation for 4D point cloud scan matching either rely on heuristic utilization of radar cross section (RCS) [huang2024less] or leverage the uncertainty inherent in point clouds derived from 3D information [xu2024modeling]. Although FMCW-based 4D ranging sensors provide Doppler velocity but still the correspondence estimation methods typically rely solely on 3D information [herraez2024radar, hexsel2022dicp, li20234d, michalczyk2022tightly, michalczyk2023multi, zhang20234dradarslam].
@@
Figure 2: The result of the proposed non-iterative scan matching method using the nyl trajectory in the NTU4DRadLM dataset [zhang2023ntu4dradlm].
@@ II Related Work
4D radar odometry and simultaneous localization and mapping (SLAM) methods have gained significant attention due to their robustness in challenging environments, such as visually degraded or bad weather conditions. Early work by doer2020ekf, doer2020radar demonstrated the feasibility of fusing 4D radar data with inertial measurements to achieve 4D radar-inertial odometry (RIO). Using extended kalman filter (EKF)-based approaches, they highlighted improvements in pose estimation through techniques such as RANSAC-based ego-velocity estimation, barometric height fusion, and online 4D radar extrinsic calibration, which removed the need for tedious pre-calibration processes. Building on this foundation, michalczyk2022tightly, michalczyk2023multi extended RIO methods, incorporating persistent landmark detection and multi-state estimation frameworks to improve performance.
More recently, as 4D radar point clouds become more dense, researchers have focused on optimization-based approaches. Unlike earlier Kalman filter-based methods, which rely on sequential state estimation, optimization-based methods aim to solve the sub-trajectory by considering the global consistency of sensor data. li20234d developed a pose graph optimization-based 4D radar SLAM framework that utilizes ego-velocity pre-integration factors to improve robustness in noisy 4D radar data. huang2024less introduced an enhanced RIO that integrates Doppler velocity and RCS information to filter noisy point cloud, adhering to the less is more principle. Additionally, nissov2024degradation tackled the challenges of LiDAR degeneracy by proposing a LiDAR-4D radar-inertial fusion, leveraging 4D radar’s robustness in the fog-filled hallway, while preserving LiDAR accuracy in well-conditioned environments. herraez2024radar presented methods for pose estimation and map creation, highlighting the benefits of velocity-aided odometry and map filtering to enhance accuracy. zhang20234dradarslam and xu2024modeling further advanced 4D radar odometry by modeling point uncertainty in polar coordinates, integrating into data association and motion estimation for improved performance in adverse conditions.
@@
4D LiDAR is an emerging technology that enables Doppler velocity measurements, offering advantages in environments with geometric degeneracy or dynamic objects. Several recent studies have explored the potential of this novel sensing modality for odometry and SLAM tasks. wu2022picking introduced the first continuous-time 4D LiDAR-only odometry method leveraging Doppler velocity measurements. Their method employs Gaussian process regression to estimate vehicle trajectories and correct motion distortion caused by scanning, significantly outperforming existing methods. Similarly, hexsel2022dicp proposed DICP, which integrates Doppler velocity measurements into the ICP framework for robust point cloud registration. By jointly optimizing Doppler and geometric objective functions, their method demonstrated improved registration accuracy and convergence speed, particularly in featureless environments such as tunnels and hallways. yoon2023need and lisus2024doppler developed correspondence-free methods that estimate 6-DOF velocity directly from Doppler information without requiring explicit scan matching and incorporate multiple sensors or complementary modules such as IMUs or gyroscopes for full motion estimation.
@@ II-C Summary and Limitations of 4D Radar / LiDAR Odometry
The 4D information of FMCW 4D LiDAR and 4D radar has leveraged its advantages in three main directions. First, Doppler velocity has been effectively incorporated into optimization frameworks to improve accuracy and robustness in challenging environments, such as repetitive geometric settings, foggy conditions, or a significant number of dynamic objects [doer2020ekf, doer2020radar, herraez2024radar, hexsel2022dicp, li20234d, michalczyk2022tightly, michalczyk2023multi, nissov2024degradation, wu2022picking, zhang20234dradarslam]. Second, scan matching methods have been developed by refining point correspondences, using RCS information [huang2024less] or the uncertainty of point clouds [xu2024modeling]. Third, direct motion estimation methods [yoon2023need, lisus2024doppler] attempt to recover full 6-DOF motion from Doppler information without relying on point correspondences.
@@ III-A Problem Definition
Mostly, the Correspondence Estimation step solely relies on the spatial information of each point cloud. Relying only on the spatial information in Correspondence Estimation makes registration prone to degradation in environments with repetitive geometric structures [hexsel2022dicp, nissov2024degradation] or when consecutive scans fail to capture the same spatial points due to its sparse and noisy characteristics [huang2024less]. To address these limitations, this study proposes Doppler Correspondence for the Correspondence Estimation, which is explained from the following in detail.
@@ III-B Geometric Derivation of Doppler Correspondence
Note that the above relationship, termed as Doppler Correspondence, is formulated such that the left-hand side, $f(p_{i},v_{p,i})$, depends only on the information from point $p_{i}$, while the right-hand side, $g(q_{j},v_{q,j})$, depends only on the information from point $q_{j}$. Furthermore, since there are no terms involving the $\Delta t$, this Doppler Correspondence is inherently invariant to translation of the sensor. Moreover, unlike traditional correspondence that relies only on spatial $(x,y,z)$ information, this method additionally utilizes Doppler information. This enables a more robust and reliable matching process in repetitive geometric environments, as it relies more on the flow of the point cloud [ding2022self] rather than the spatial information from the environments.
@@ III-C Kinematic Derivation of Doppler Correspondence
Alternative to the previous geometric derivation of Doppler Correspondence, this section utilizes manipulation of simple matrices and vectors. To establish Doppler Correspondence, we start with the kinematic relationship from  [ding2022self, zhou2023self]:
@@ III-D Outlier Rejection
This occurs because $g(q_{j},v_{q,j})$ is identical to $g(q^{\prime}_{j},v_{q^{\prime},j})$, as the range and radial velocity of each point is the same ($r_{q_{j}}=r_{q^{\prime}_{j}}\text{ and }v_{q_{j}}=v_{q^{\prime}_{j}}$). Such ambiguity can result in incorrect correspondence estimation. To address this issue, it is assumed that correctly matched points $p_{i}$ and $q_{j}$ lie within similar spatial regions between scans. This assumption enables the use of a distance-based rejection mechanism to filter out mismatched candidates, as shown in Fig. 4(b). Similar to the maximum correspondence distance used in [segal2009generalized], a threshold is applied to eliminate unlikely matches. Specifically, both the spatial distance $d_{\text{spatial}}=\|p_{i}-q_{j}\|_{2}$ and the Doppler distance $d_{\text{doppler}}=|f(p_{i},v_{p,i})-g(q_{j},v_{q,j})|$ should satisfy:
@@ III-E Non-iterative Scan Matching Algorithm
The process begins by receiving two consecutive scans, $\mathcal{P}$ and $\mathcal{Q}$, and calculating $f(p_{i},v_{p,i})\text{ and }g(q_{j},v_{q,j})$. FindClosest is then performed to estimate correspondences. Unlike the conventional ICP explained in Algorithm 2, which relies on 3D-based correspondence, our method operates on 1D-based correspondence. This could reduce computational complexity, even though $f(p_{i},v_{p,i})$ and $g(q_{j},v_{q,j})$ need to be calculated. While this approach may be slightly less accurate than 3D-based methods due to less information, it offers faster computations. Furthermore, unlike the closest point correspondence methods [besl1992method, hexsel2022dicp, rusinkiewicz2001efficient, segal2009generalized], which iteratively refine correspondences through multiple iterations, Algorithm 1 performs a direct correspondence estimation in a single step. This eliminates the iterative search process, significantly reducing the computational load, as illustrated in Fig. 1. Next section, the feasibility of Doppler Correspondence is validated by comparing the proposed method against ICP-based approaches.
@@ IV-A1 Sparse 4D Radar
The Campus dataset [li20234d] is utilized to evaluate sparse 4D radar point clouds. The data were collected using a ZF FRGen21 4D radar sensor mounted on the front bumper of a vehicle. Each 4D radar scan generates a sparse and noisy point cloud, containing approximately 400 to 1,400 points every 60 ms. The Campus dataset encompasses outdoor environments featuring trees, vehicles, and pedestrians. These diverse settings challenge methods to operate effectively under low-density radar point cloud conditions.
@@ IV-A2 Semi-Dense 4D Radar
The Loop 1 dataset from the NTU4DRadLM dataset [zhang2023ntu4dradlm] is used to evaluate semi-dense 4D radar. The data were collected using a car platform with an Oculii Eagle 4D radar, traversing the NTU campus main roads, resulting in a 6.95 km trajectory. The car moved at an average speed of 25–30 km/h, and the 4D radar scans exhibit semi-dense point clouds with approximately 1,000 to 2,000 points per 80ms. The environment features a mix of structured and semi-structured regions.
@@ IV-A3 Dense 4D LiDAR
The Straight Wall, Curved Wall, and Baker-Barry Tunnel datasets from [hexsel2022dicp] are utilized to evaluate dense 4D LiDAR. The Straight Wall and Curved Wall datasets, generated in the CARLA simulator, simulate featureless environments by placing large parallel walls along the trajectory. These datasets provide an effective platform for evaluating odometry performance, particularly in environments with limited geometric features and repetitive structures. The Baker-Barry Tunnel dataset was collected using an Aeva Aeries I FMCW 4D LiDAR sensor. The point clouds from 4D LiDAR are significantly denser than 4D radar data, containing approximately 40,000 to 120,000 points per 100ms, depending on the sequence.
@@ IV-B Experiment Setup
The proposed method is evaluated by comparing its performance with established methods: point-to-point ICP and DICP [hexsel2022dicp]. DICP is a novel algorithm that integrates Doppler velocity into the Optimization in the ICP framework. It enhances scan matching by jointly optimizing both the closest point term and additional constraints derived from the Doppler velocity of each point. These methods are tested in two conditions: No Seed and With Seed. In the No Seed category, no initial pose estimate is provided, whereas the With Seed category uses the estimated pose from the previous scan pair under a constant-velocity motion model assumption [vizzo2023kiss].
@@
Semi-Dense 4D Radar Loop1_0 (With Seed)DICP 0.821 / 0.627 1.201 / 1.110 1.434 / 1.606 1.600 / 2.084 1.771 / 2.559 1.949 / 3.024 4.7
@@
Dense 4D LiDAR Straight Wall (With Seed)DICP 0.050 / 0.028 0.111 / 0.059 0.128 / 0.073 0.154 / 0.086 0.258 / 0.103 0.303 / 0.112 29.1
@@
Baker-Barry Tunnel (Empty) (With Seed)DICP 1.028 / 0.289 1.652 / 0.490 2.515 / 0.744 3.313 / 0.963 3.964 / 1.140 4.794 / 1.342 20.8
@@
Baker-Barry Tunnel (Vehicles) (With Seed)DICP 0.778 / 0.179 1.435 / 0.339 2.110 / 0.497 2.525 / 0.621 3.027 / 0.732 3.479 / 0.833 20.1
@@ IV-C Evaluation Metrics
The performance is evaluated using the relative error (RE) from [zhang2018tutorial]. This metric standardizes trajectory estimation accuracy, allowing fair comparisons across different methods. The evaluation measures odometry accuracy through segment-wise error analysis. This approach provides detailed insights into the local and global performance of the trajectory estimation. By using shorter sub-trajectories, RE captures local consistency and short-term accuracy, while longer sub-trajectories emphasize global consistency and long-term accuracy. This dual capability makes RE a versatile and informative metric for trajectory analysis.
@@ IV-D Experiment Results
The performance evaluation on sparse 4D radar data compares the proposed method against the standard point-to-point ICP approach, while the open-source implementation of DICP is excluded due to its failure to produce reliable results on sparse 4D radar data. Under these conditions, closest point correspondence, which relies on the spatial overlap of consecutive scans, becomes unreliable [huang2024less]. By contrast, the proposed correspondence leverages the flow of points [ding2022self] rather than explicit spatial positions, offering robustness against sparse and noisy data. As shown in Table I, Figs. 5(a) and (b), and Fig. 6(a), the proposed method attains substantially higher accuracy than that of point-to-point ICP. Moreover, unlike point-to-point ICP, which requires repeated correspondence estimation in 3D space, our approach simplifies the process to a single-step and 1D correspondence estimation.